A plane diffraction grating is illuminated by a source that emits two spectral lines of wavelengths 420nm and 600nm. Show that the 3rd order of the line of one of these wavelengths is diffracted through a greater a greater angle than the 4th order of the other wavelength.


1
Expert's answer
2021-02-17T12:07:12-0500

Answer

"\lambda_1=420nm\\\lambda_2=600nm"


n1=3

n2=4

Using bragg equation

"2dsin\theta=n\lambda"

Here interplaner distance and wavelength is different.

So angle relationship

"\frac{sin\theta_1}{sin\theta_2 }=\frac{n_1\lambda_1}{n_2 \lambda_2 }=\frac{3\times420}{4\times600}"

"{sin\theta_1}=0.525\times sin\theta_2"

Which means "\theta_1" Is lesser.